# Design principles of the cytotoxic CD8+ T-cell response

### v2

@@ Results
Varying these baseline rates produced strikingly different effector trajectories (Fig. 2). Promoting activation/proliferation ($g_{0,N\to N^{*}},g_{0,E\to 2E}\,\uparrow$), or lowering the contraction baseline rate ($g_{0,E\to\emptyset}\,\downarrow$) generated a rapid, high-amplitude effector burst that cleared the pathogen quickly, after which effector numbers decay (Fig. 2B, C). In contrast, slowing activation and division ($g_{0,N\to N^{*}},g_{0,E\to 2E}\,\downarrow$), or accelerating contraction ($g_{0,E\to\emptyset}\,\uparrow$), yielded sluggish effector accumulation, delayed clearance, and in extreme cases pathogen persistence, accompanied by a prolonged plateau of few effectors. Interestingly, varying the baseline rate of differentiation $g_{0,N^{*}\to E}$ had a weaker impact on the overall effector profile and pathogen clearance (Fig. 2B, C), suggesting relevance more downstream in the response program.
Because effector and pathogen abundances determine the levels of antigen and cumulative harm-associated cytokine signals, these baseline choices not only shape the immediate effector response but also alter the available antigen and cytokine signals, which in turn impact the T-cell response through feedback regulation (SI Appendix, Fig. S2).
@@
Previous research has identified a trade-off between maximizing infection clearance and minimizing response toxicity (immunopathology) in the immune response [61, 28, 29]. Consistently, by systematically sampling the space of designs (SI Appendix), we found this trade-off in the form of a Pareto front: plotting infection harm $H_{\text{inf.}}(\theta)$ (i.e., infected cells killed by pathogen or by effector response) versus the toxicity $H_{\text{toxic.}}(\theta)$ (i.e., bystander cells killed by effector response) revealed that decreasing $H_{\text{inf.}}$ comes at the cost of increasing $H_{\text{toxic.}}$ (Fig. 3A). The extent of this trade-off depends on pathogen characteristics. This trade-off is stronger for infections with a higher intra-host reproductive number $R_{0}$ or lower immunogenicity $\kappa$. Interestingly, for scenarios with lower $R_{0}$, many of the hard-wired response programs without signal feedback ($\vec{\psi}_{i}=0$) lie along the Pareto front, and perform near-optimally. For aggressive high-$R_{0}$ threats—be it acute infections (high $\kappa$), or highly metastatic cancers (mid to low $\kappa$)—these same hard-wired designs fall far inside the sub-optimal region. Proportionate control in these cases requires designs with appropriate sensitivities to antigen and cytokine/harm signals ($\vec{\psi}\neq 0$) to minimize infection harm while limiting immunopathology.
@@
Figure 4: Pareto-optimal designs balancing harm from acute infections and autoimmunity. (A) Bi-objective optimization to minimize the weighted sum of harms associated with T-cell regulatory designs $\vec{\theta}$ across two infection archetypes ($A,B$) yields a Pareto front (magenta curve); eq. 10. Points on the curve are Pareto-optimal designs, parameterized by the weight $\lambda\in[0,1]$, quantifying the emphasis on archetype B; moving left to right corresponds to decreasing $\lambda$ (less emphasis on B). (B) The density plot shows harms incurred by $4.8\times 10^{6}$ uniformly sampled designs in their response to archetype I (non-replicating, self-like/ auto-immune inducing or allergenic antigens; $R_{0}=1$) versus archetype II (immunogenic infection; high $\kappa$), as defined in Fig. 3B. Designs are sampled on a uniform Cartesian grid of sensitivities $\psi_{\sigma}\in[-3,3]$ and baseline rates $\ell_{i}\in[-3,3]$ (as in Fig. 3A). For many values of $\lambda$, we minimize the objective in (A), and trace out the Pareto front (magenta). For each point $\lambda$ on the front, we identify the ensemble of designs closest to the Pareto front ($\Theta(\lambda)$), where a design belongs to the near-Pareto ensemble at $\lambda$ if its incurred harm against archetypes I and II simultaneously lie within 5% of the $\lambda$-Pareto-optimal harm of the respective archetypes. The black square indicates the elbow of the Pareto curve $\lambda^{*}$: the point where further reductions in archetype-I harm incur larger increases in archetype-II harm (see SI Appendix). The average design parameter values over the near-elbow ensemble are $\langle\vec{\theta}\rangle_{\Theta(\lambda^{*})}=(2.89,1.67,-0.67,-0.22,-1.94,0.94)$. (C) Decomposition of total harm along the Pareto front in (B) into infection-related harm and toxicity (eq. 9). The black dashed line indicates the position of the elbow from (B). (D) Solid lines show parameter magnitudes for the signal sensitivities $\psi$ (left) and the baseline rates $g_{0}(\ell)$ (right), each scaled to variable bounds $(\psi^{\max}=3,\ \ell^{\max}=3)$, as a function of harm incurred under archetype II along the Pareto curve (parameterized by $\lambda$). Shaded envelopes denote the standard deviation of the near-Pareto ensemble of designs $\Theta(\lambda),\penalty 10000\ \lambda\in[0,1]$ traced out by the Pareto front. (E) For each archetype in Fig. 3B, total harm is shown for the archetype-specific minimizer $\vec{\theta}^{(\mathrm{min})}$ (circle) and for the average over near-elbow ensemble in (B). See SI Appendix, Figs. S6, S7 for Pareto fronts formed by other pairs of archetypes. All simulation parameters other than the varied $K_{I}$, $b_{I}$, and $\vec{\theta}$ match Fig. 1.
@@
Next, we explored regulatory design space by evaluating responses of different designs against diverse infection scenarios, sampled on a grid of immunogenicities ($\kappa$) and reproduction numbers $(R_{0})$ (Fig. 3B). For each infection scenario $(\kappa,R_{0})$, we identified the ensemble of near-optimal designs, $\Theta(\kappa,R_{0})$. A design $\vec{\theta}$ belongs to this ensemble if its incurred total harm (eq. 9) is within $10^{-2}\times S_{\text{max}}$ of the minimum realized total harm from all sampled designs for that infection. Finally, we clustered infections based on their ensemble-averaged design parameters $\langle\vec{\theta}\rangle_{\Theta(\kappa,R_{0})}$ (SI Appendix, Fig. S3). Four coherent design classes emerged, loosely corresponding to intuitive pathogen “archetypes” (Fig. 3B): I—non-replicating, self-like/auto-immune inducing or allergenic antigens ($R_{0}=1$); II—highly immunogenic bacteria and viruses (high $\kappa$ or high $R_{0}$); III—intermediate replication and detectability, typifying mixed-strategy intracellular pathogens (mid-high $R_{0}$, mid-high $\kappa$); IV—difficult to detect threats such as growing tumors (low $\kappa$, broad range of $R_{0}$). The Pareto trade-off between infection harm and toxicity differs across these archetypes: designs optimized for archetypes I-III nearly eliminate infection damage with limited immunopathology, while designs in archetype IV show a strong trade-off (Fig. 3C).
@@
Design parameters of effector programs minimizing harm vary across archetypes (Fig. 3D). For archetype I—non-replicating antigens—the optimal design is largely pre-programmed to suppress the immune response: signal sensitivities $\vec{\psi}$ are close to 0, and the baseline rates are tuned to slow activation, proliferation, and differentiation, and enforce rapid effector deletion. This damped response completely avoids toxicity when the threat does not grow so infection harm is determined by the initial exposure and cannot be decreased. Across all other archetypes a common pattern emerges—the response-derived cytokine signal $\sigma_{\text{res}}$ acts as an anti-inflammatory brake ($\psi_{\text{res}}<0$), with its magnitude tuned to context. Moreover, antigen sensitivity $\psi_{\text{Ag}}$ grows with immunogenicity: the effector program optimized for highly immunogenic pathogens in archetype II shows the largest $\psi_{\text{Ag}}$, whereas the programs for the poorly immunogenic archetype IV show diminished antigen sensitivity. Compared to archetype II, the baseline propensities $\ell$’s (or equivalently $g_{0}$’s) for archetype III accelerate effector contraction ($g_{0,E\to\emptyset}\,\uparrow$). For the weakly immunogenic archetype IV, near-optimal designs distinctively tune both the baseline rates $g_{0}$’s and the infection-signal sensitivity $\psi_{\text{inf}}$ to balance clearance against toxicity and minimize harm (SI Appendix, Figs. S4, S5).
@@
Although the relative evolutionary pressures exerted by different threats are unknown, we can still pose a principled multi-objective problem for the effector program, i.e., find a design $\vec{\theta}$ that minimizes the weighted sum of harms across pathogen archetypes:
$\vec{\theta}^{\prime}(\{\lambda_{i}\})=\underset{\vec{\theta}}{\arg\min}\sum_{i:\text{archetypes}}\lambda_{i}H^{(i)}(\vec{\theta}),$ (10)
where $\lambda_{i}\in[0,1]$ with the constraint $\sum_{i}\lambda_{i}=1$. The weights $\{\lambda_{i}\}$ determine the relative importance of each archetype in determining the optimal design $\vec{\theta}^{\prime}$. The solution to this family of problems forms a simplex over which $\lambda_{i}^{\prime}s$ vary (Fig. 4A; SI Appendix). This approach accommodates uncertainty about the empirical prevalence of each type of infection and the preferences/contingencies of evolution.
@@
Even though optimization can be done simultaneously over all four archetypes, we proceed by examining pairwise trade-offs that generate a one dimensional Pareto front, parameterized by $\lambda$ (Fig. 4A). A biologically salient case contrasts archetype I (non-replicating antigens) with archetype II (immunogenic pathogens). Non-replicating antigens plausibly arise from unprogrammed cell death events or external trauma inducing inflammation and may be implicated in autoimmune flares. This bi-objective harm minimization aims to find designs balancing effective clearance of immunogenic viral or bacterial infections and suppression of autoimmune responses; see SI Appendix, Figs. S6, S7, for Pareto fronts formed by optimizing other pairs of clusters.
@@
Fig. 4B shows harm from archetype I plotted against that of archetype II for a broad range of sampled designs and highlights the Pareto front (magenta). Designs close to the Pareto front achieve near-best total protection for a given compromise weight $\lambda$; moving to the right along the Pareto front shifts the emphasis from clearance of immunogenic pathogens ($\lambda\uparrow$) to tolerance ($\lambda\downarrow$). Decomposing total harm into infection-related harm and toxicity (eq. 9) shows that along this Pareto front it is relatively easy to reduce toxicity to non-replicating antigens without substantially worsening outcomes against immunogenic infections (Fig. 4C).
@@
Figure 5: Elbow-optimal designs recapitulating macro-dynamics of T-cell expansion. (A,B) Fold expansion (i.e, peak clone size divided by the initial precursor number) of antigen-specific CD8+ T-cells in mice plotted against (A) the initial precursor number (data from [9]), and (B) the relative pMHC concentration required for half-maximal activation EC50 (data from [62]); in our model, EC50 is proxied by $K_{I}$. (C,D) Simulated mean fold expansion for designs belonging to the near-elbow ensemble from Fig. 4B ($\Theta(\lambda^{*})$) when their sensitivity to the infection-harm signal ($\psi_{\text{inf}}$) is varied (colors), shown as a function of (C) the initial precursor T-cell number, and (D) the immunogenicity of the antigen $\kappa$. For each design in the near-elbow ensemble identified in Fig. 4, $\vec{\theta}\in\Theta(\lambda^{*})$, we change $\psi_{\text{inf}}\to c\times\psi_{\text{inf}}$ for constants $c\in\{0,0.1,1.0\}$ (colors), and simulate responses against the broad sample of infections (Fig. 3B). Trend lines are obtained by averaging over infections and designs in the near-elbow ensemble (SI Appendix, Figs. S8, S9). Elbow designs with reduced sensitivity to infection-derived harm qualitatively match the experimental macro-dynamics of T-cell expansion. Unless otherwise indicated, simulation parameters: $S_{\text{max}}=10^{7},K_{S}=S_{\text{max}},K_{H}=10^{-3}S_{\text{max}},d_{I}=0.5\penalty 10000\ \text{day}^{-1},I_{0}=10^{-4}S_{\text{max}}$.
Table 1: Endogenous design of CD8+ T-cell response.
@@
Next, we evaluated immune designs near the Pareto curve (Fig. 4D; SI Appendix). For each point $\lambda$, we identified an ensemble of designs near the Pareto front, $\Theta(\lambda)$. For a design to belong to the near-Pareto ensemble, its incurred harm against archetypes I and II must simultaneously lie within 5% of the $\lambda$-Pareto-optimal harm of those respective archetypes. The elbow of the Pareto curve represents the point on the curve, $\lambda^{*}$, at which further decreases in harm from archetype I will incur disproportionally large increases in harm from archetype II (see SI Appendix). The ensemble of designs near the elbow $\Theta(\lambda^{*})$ encodes a compelling logic: sensitivity from antigen and infection cues promotes effector production ($\psi_{\text{Ag}},\psi_{I}>0$), whereas response-derived cues act like a brake with a predominantly anti-inflammatory effect on the immune response ($\psi_{\text{res}}<0$). Sensitivity to these signals diminishes as $\lambda$ decreases (i.e., as avoidance of autoimmunity is prioritized); see Fig. 4D. Baseline rates ($g_{0}$’s) of the designs at the elbow favor moderate to slow activation and proliferation of effectors, slow differentiation to effector state, and moderate contraction (Fig. 4D). As $\lambda$ increases (leftward; clearance weighed more), all baseline rates shift toward promoting more effector cells. As $\lambda$ decreases below the elbow (rightward), the already slow differentiation baseline rate falls toward zero, while activation/proliferation and contraction are tuned to suppress effector production, reflecting designs that minimize collateral damage even at the expense of incomplete clearance. Notably, despite being optimized only for archetypes I and II, designs in the near-elbow ensemble $\Theta(\lambda^{*})$ perform competitively for other archetypes, incurring harm close to the archetype-specific optimal designs (Fig. 4E; SI Appendix, Figs. S6, S7).
@@
Figure 6: Engineering immune response for cancer immunotherapy. (A) The schematic shows the potential clearance of replicating cancer tumor cells by T-cells (e.g., infused during immunotherapy), and the potential collateral damage of therapy on susceptible tissue cells. The accessible levers in an immunotherapy include the dosage and antigenicity of the infused cells. Antigenicity is a property of the interaction between a specific TCR-antigen pair, and it can be increased by either changing the antigen or the TCR. Given a specific cognate T-cell, immunogenicity is the ratio of the antigenicity of a cancer cell to that of a healthy cell. (B) Averaged over the tuned near-elbow ensemble $\tilde{\Theta}(\lambda^{*})$, the heatmap displays harm as a function of cancer antigenicity and dosage of administered T-cells. Yellow points correspond to the optimal dosage of T-cells for a given immunogenicity/antigenicity. (C) Total harm, harm from cancer and toxicity as functions of cancer immunogenicity ($\kappa$) are shown at the optimal T-cell dosage (yellow circles in B). (D) Clearance response—defined as the derivative of negative harm (i.e., protection) with respect to each design parameter $\theta_{i}$, $\partial H(\vec{\theta})/\partial{\theta_{i}}$ for $\vec{\theta}\in\tilde{\Theta}(\lambda^{*})$—averaged over the tuned near-elbow designs $\tilde{\Theta}(\lambda^{*})$, is shown for different levels of tumor immunogenicity $\kappa$; Left: $\vec{\psi}$ parameters; right: $\vec{g}_{0}$ parameters (colors). Shaded envelopes denote the standard deviation of outcomes over $\theta\in\tilde{\Theta}(\lambda^{*})$. For this analysis, we assume that T-cells are administered at high dosage ($N_{\text{lin.}}=10^{-2}S_{\text{max}}$). Simulation parameters: $S_{\text{max}}=10^{7},K_{S}=S_{\text{max}},K_{H}=10^{-3}S_{\text{max}},d_{I}=0.0\penalty 10000\ \text{day}^{-1},I_{0}=10^{-1}S_{\text{max}},b_{I}=0.1\penalty 10000\ \text{day}^{-1}$.
@@
The choice of design impacts the dynamics of the immune response, which we can compare to experimental data. Prior experiments in mice [62, 9] indicate that peak fold-expansion of T-cell populations in response to a pathogen scales inversely with the initial number of cognate T-cells, and with the pMHC concentration for half-maximal response (proxied by $K_{I}$ in our model); see Fig. 5A, B. This pattern was recapitulated by explicitly modeling competition for antigens in polyclonal T-cell responses [65]. Near-elbow designs $\vec{\Theta}(\lambda^{*})$ (Fig. 4) reproduce this inverse dependence on initial clone size but underestimate the steepness of the trend with respect to immunogenicity ($\kappa$) (Fig. 5C, D). Reducing the sensitivity to infection harm $\psi_{\text{inf}}$ of these designs by an order of magnitude restores the observed dynamic range (Fig. 5D) while having only a minor effect on total harm, and hence, our optimization objective (Fig. S9). Interestingly, adjusting the other sensitivities ($\psi_{\text{Ag}},\psi_{\text{res}}$), or the baseline rates $g_{0}$’s does not reproduce the dynamic range observed in the experiments (Figs. S8, S9).

### v3

@@ III Results
Varying these baseline rates produced strikingly different effector trajectories (Fig. 2). Promoting activation/proliferation ($g_{0,N\to N^{*}},g_{0,E\to 2E}\,\uparrow$), or lowering the contraction baseline rate ($g_{0,E\to\emptyset}\,\downarrow$) generated a rapid, high-amplitude effector burst that cleared the pathogen quickly, after which effector numbers decay (Fig. 2B, C). In contrast, slowing activation and division ($g_{0,N\to N^{*}},g_{0,E\to 2E}\,\downarrow$), or accelerating contraction ($g_{0,E\to\emptyset}\,\uparrow$), yielded sluggish effector accumulation, delayed clearance, and in extreme cases pathogen persistence, accompanied by a prolonged plateau of few effectors. Interestingly, varying the baseline rate of differentiation $g_{0,N^{*}\to E}$ had a weaker impact on the overall effector profile and pathogen clearance (Fig. 2B, C), suggesting its relevance may lie elsewhere in the response program.
Because effector and pathogen abundances determine the levels of antigen and cumulative harm-associated (cytokine) signals, these baseline choices not only shape the immediate effector response but also alter the available antigen and cytokine signals, which in turn impact the T-cell response through feedback regulation (SI Appendix, Fig. S2).
@@
Previous research has identified a trade-off between maximizing infection clearance and minimizing response toxicity (immunopathology) in the immune response [59, 28, 29]. Consistently, by systematically sampling the space of designs (SI Appendix), we found the same trade-off: plotting infection harm $H_{\text{inf.}}(\theta)$ (i.e., infected cells killed by pathogen or by effector response) versus the toxicity $H_{\text{toxic.}}(\theta)$ (i.e., bystander cells killed by effector response) revealed that decreasing $H_{\text{inf.}}$ comes at the cost of increasing $H_{\text{toxic.}}$ (Fig. 3A). The extent of this trade-off depends on pathogen characteristics. This trade-off is stronger for infections with a higher intra-host reproductive number $R_{0}$ or lower immunogenicity $\kappa$. Interestingly, for scenarios with low $R_{0}$, many of the hard-wired response programs without signal feedback ($\vec{\psi}_{i}=0$; black dots) perform near-optimally, laying close to the Pareto front. Here, the Pareto fronts at the boundaries of these graphs indicate the set of strategies for which neither harm nor toxicity can be reduced without increasing the other, representing the best achievable balance between these competing objectives (see SI Appendix). For aggressive high-$R_{0}$ threats—be it acute infections (high $\kappa$), or highly metastatic cancers (mid to low $\kappa$)—these same hard-wired designs fall far from the Pareto front and well inside the sub-optimal region. Proportionate control in these cases requires designs with appropriate sensitivities to antigen and cytokine/harm signals ($\vec{\psi}\neq 0$) to minimize infection harm while limiting immunopathology.
@@
Figure 4: Pareto-optimal designs balancing harm from acute infections and autoimmunity. (A) Bi-objective optimization to minimize the weighted sum of harms associated with T-cell regulatory designs $\vec{\theta}$ across two pathogen archetypes ($A,B$) yields a Pareto front (magenta curve); eq. 8. Points on the curve are Pareto-optimal designs, parameterized by the weight $\lambda\in[0,1]$, quantifying the emphasis on archetype A; moving left to right corresponds to decreasing $\lambda$ (less emphasis on archetype A). (B) The density plot shows harms (in units of the maximum number of susceptible cells $S_{\text{max}})$ incurred by $4.8\times 10^{6}$ uniformly sampled designs in their response to archetype I (immunogenic infection; high $\kappa$) versus archetype II (non-replicating, self-like/ auto-immune inducing or allergenic antigens; $R_{0}=1$), as defined in Fig. 3B. Designs are sampled on a uniform Cartesian grid of sensitivities $\psi_{\sigma}\in[-3,3]$ and baseline rates $\ell_{i}\in[-3,3]$ (as in Fig. 3A). For many values of $\lambda$, we minimize the objective in (A), and trace out the Pareto front (magenta). For each point $\lambda$ on the front, we identify the ensemble of designs closest to the Pareto front ($\Theta(\lambda)$), where a design belongs to the near-Pareto ensemble at $\lambda$ if its incurred harm against archetypes I and II simultaneously lie within 5% of the $\lambda$-Pareto-optimal harm of the respective archetypes. The black square indicates the elbow of the Pareto curve $\lambda^{*}$: the point where further reductions in archetype-II harm incur larger increases in archetype-I harm (SI Appendix). (C) Decomposition of total harm along the Pareto front in (B) into infection-related harm and toxicity (eq. 7). The black dashed line indicates the position of the elbow from (B). (D) The magnitudes of the signal sensitivities $\psi$ (left) and the baseline rates $g_{0}(\ell)$ (right), each scaled to variable bounds $(\psi^{\max}=3,\ \ell^{\max}=3)$, are shown as a function of harm incurred under archetype I along the Pareto curve (parameterized by $\lambda$). Solid and shaded envelopes show the mean and standard deviation of the design parameters, respectively, across the near-Pareto ensemble of designs $\Theta(\lambda),\penalty 10000\ \lambda\in[0,1]$ traced out by the Pareto front. (E) For each archetype in Fig. 3B, total harm is shown for the archetype-specific minimizer $\vec{\theta}^{(\mathrm{min})}$ (circle) and for the average over the near-elbow ensemble in (B). See SI Appendix, Fig. S6 for Pareto fronts formed by other pairs of archetypes. All simulation parameters other than the varied $K_{I}$, $b_{I}$, and $\vec{\theta}$ match Fig. 1.
@@
Next, we explored regulatory design space by evaluating responses of different designs against diverse infection scenarios, sampled on a grid of immunogenicities ($\kappa$) and reproduction numbers $(R_{0})$ (Fig. 3B). For each infection scenario $(\kappa,R_{0})$, we identified the ensemble of near-optimal designs, $\Theta(\kappa,R_{0})$. A design $\vec{\theta}$ belongs to this ensemble if its incurred total harm (eq. 7) is within $10^{-2}\times S_{\text{max}}$ of the minimum realized total harm from all sampled designs for that infection; note that the maximum number of susceptible cells $S_{\text{max}}$ provides a convenient reference scale against which to quantify harm (cell death). Finally, we clustered infections based on their ensemble-averaged design parameters $\langle\vec{\theta}\rangle_{\Theta(\kappa,R_{0})}$ (SI Appendix, Fig. S3). Three coherent design classes emerged, loosely corresponding to intuitive pathogen “archetypes” (Fig. 3B): I— highly immunogenic bacteria and viruses (high $\kappa$ or high $R_{0}$), associated with type-I immunity [60]; II— non-replicating, self-like/auto-immune inducing or allergenic antigens ($R_{0}=1$), in part associated with type-II immunity [60]; III— mid-range threats (low-medium $\kappa$ and $R_{0}$) including difficult to detect pathogens and slowly growing tumors. The balance between minimizing toxicity and clearing infection (i.e., reducing harm) differs across these archetypes, resulting in different design strategies: signal-sensitive designs optimized for archetypes I and III prioritize pathogen clearance at the cost of increased immunopathology, with the severity of the trade-off varying by scenario, whereas the largely pre-programmed designs optimized for archetype II suppress toxicity to avoid autoimmunity (Fig. 3C).
@@
Design parameters of effector programs minimizing harm vary across archetypes (Fig. 3D). For archetype II (non-replicating antigens) the optimal design is pre-programmed suppression of the immune response: signal sensitivities $\vec{\psi}$ are close to 0, and the baseline rates are tuned to slow activation, proliferation, and differentiation, and enforce rapid effector deletion. This damped response completely avoids toxicity when the threat does not grow so infection harm is determined by the initial exposure and cannot be decreased. Across the other two archetypes a common pattern emerges: the response-derived cytokine signal $\sigma_{\text{res}}$ acts as an anti-inflammatory brake ($\psi_{\text{res}}<0$), with its magnitude tuned to context. Moreover, antigen sensitivity $\psi_{\rm Ag}$ grows with immunogenicity, with the designs optimized for archetype I (strong immunogenicity) showing the largest sensitivity to antigens, while those for archetype II (non-immunogenic) show no antigen sensitivity, and are largely hard-wired. Overall, across all archetypes, near-optimal designs distinctively tune both the baseline rates $g_{0}$’s and the infection-signal sensitivity $\psi_{\text{inf}}$ to balance clearance against toxicity and minimize harm (SI Appendix, Figs. S4, S5).
@@
Although the relative evolutionary pressures exerted by different threats are unknown, we can still pose a principled multi-objective problem for the effector program, i.e., find a design $\vec{\theta}_{\text{min}}$ (eq. 6) that minimizes the weighted sum of harms across a desired set of pathogen archetypes:
$\vec{\theta}_{\text{min}}=\underset{\vec{\theta}}{\arg\min}\sum_{i:\text{archetypes}}\lambda_{i}H^{(i)}(\vec{\theta}),$ (8)
where $\lambda_{i}\in[0,1]$ with the constraint $\sum_{i}\lambda_{i}=1$. The weight $\lambda_{i}$ specifies how strongly we prioritize minimizing harm associated with the $i^{th}$ pathogen archetype (i.e., an objective), when identifying the optimal design $\vec{\theta}_{\text{min}}$. For $m$ archetypes, the solution to this family of problems maps an m1m-1-dimensional simplex of $\vec{\lambda}=(\lambda_{1},\dots,\lambda_{m})$ values to a set of designs that achieve the best possible trade-off between the different archetypes (Fig. 4A; SI Appendix). When optimizing performance against one archetype necessarily worsens performance against another (i.e., a tradeoff), these optimal designs form a Pareto front: a set of solutions for which no objective can be improved without degrading at least one other (Fig. 4A). In contrast, if there are no such conflicts among archetypes, the Pareto front collapses to a single design that is simultaneously optimal for all objectives. This framework naturally accommodates uncertainty in the prevalence of pathogenic challenges and in the effective “preferences” imposed by evolution or environment, in finding the optimal solution. Similar approaches have been successfully applied in other biological systems [61, 62].
@@
Even though optimization can be done simultaneously over all three archetypes, we proceed by examining pairwise trade-offs that generate a one dimensional Pareto front, parameterized by a single scalar $\lambda$ (Fig. 4A). A biologically salient case contrasts archetype I (immunogenic pathogens) with archetype II (non-replicating antigens). Non-replicating antigens plausibly arise from un-programmed cell deaths or external trauma inducing inflammation and may be implicated in autoimmune flares. This bi-objective harm minimization aims to find designs balancing effective clearance of immunogenic viral or bacterial infections and suppression of autoimmune responses; see SI Appendix, Fig. S6 for Pareto fronts formed by optimizing other pairs of clusters.
@@
Fig. 4B shows harm from archetype II plotted against that of archetype I for a broad range of sampled designs and highlights the Pareto front (magenta). Designs close to the Pareto front achieve near-best total protection for a given compromise weight $\lambda$; moving to the right along the Pareto front shifts the emphasis from clearance of immunogenic pathogens ($\lambda\uparrow$) to tolerance ($\lambda\downarrow$). Decomposing total harm into infection-related harm and toxicity (eq. 7) shows that along this Pareto front it is relatively easy to reduce toxicity to non-replicating antigens without substantially worsening outcomes against immunogenic infections; this is demonstrated in Fig. 4C as substantial decreases in toxicity are achieved with only minor changes in infection harm.
@@
Figure 5: Elbow-optimal designs recapitulating macro-dynamics of T-cell expansion. (A,B) In maroon (left vertical axis), fold expansion (i.e, peak clone size divided by the initial precursor number) of antigen-specific CD8+ T-cells in mice plotted against (A) the initial precursor number (data from [9]), and (B) the relative pMHC concentration required for half-maximal activation EC50 (data from [63]); in our model, EC50 is proxied by $K_{I}$. In blue (right vertical axis), simulated mean fold expansion for designs belonging to the near-elbow ensemble from Fig. 4B ($\Theta(\lambda^{*})$) shown as a function of the initial precursor T-cell number (A), and the immunogenicity of the antigen $\kappa$ (B). Their sensitivity to the infection-harm signal ($\psi_{\text{inf}}$) is varied (shades of blue). For each design in the near-elbow ensemble identified in Fig. 4, $\vec{\theta}\in\Theta(\lambda^{*})$, we change $\psi_{\text{inf}}\to c\times\psi_{\text{inf}}$ for constants $c\in\{0,0.1,1.0\}$, and simulate responses against the broad sample of infections (Fig. 3B). Trend lines are obtained by averaging over infections and designs in the near-elbow ensemble (SI Appendix, Figs. S7, S8). Elbow designs with reduced sensitivity to infection-derived harm qualitatively match the experimental macro-dynamics of T-cell expansion. Unless otherwise indicated, simulation parameters: $S_{\text{max}}=10^{7},K_{S}=S_{\text{max}},K_{H}=10^{-3}S_{\text{max}},d_{I}=0.5\penalty 10000\ \text{day}^{-1},I_{0}=10^{-4}S_{\text{max}}$.
@@
Table 1: Endogenous design of CD8+ T-cell response.
Next, we evaluated immune designs near the Pareto front (Fig. 4D; SI Appendix). For each point $\lambda$, we identified an ensemble of designs near the Pareto front, $\Theta(\lambda)$. For a design to belong to the near-Pareto ensemble, its incurred harm against archetypes I and II must simultaneously lie within 5% of the Pareto optimal harm of those respective archetypes at the same compromise value $\lambda$. The elbow of the Pareto front represents the point on the curve, $\lambda^{*}$, at which further decreases in harm from archetype II will incur disproportionally large increases in harm from archetype I (Fig. 4B; SI Appendix). The ensemble of designs near the elbow $\Theta(\lambda^{*})$ encodes a compelling logic: sensitivity from antigen and infection cues promotes effector production ($\psi_{\text{Ag}},\psi_{I}>0$), whereas response-derived cues act like a brake with a predominantly anti-inflammatory effect on the immune response ($\psi_{\text{res}}<0$); Fig. 4D. Sensitivity to these signals diminishes as $\lambda$ decreases (i.e., as avoidance of autoimmunity is prioritized); see Fig. 4D. Baseline rates ($g_{0}$’s) of the designs at the elbow favor moderate to slow activation and proliferation of effectors, slow differentiation to effector state, and moderate contraction (Fig. 4D). As $\lambda$ increases (leftward; clearance weighes more), all baseline rates shift toward promoting more effector cells. As $\lambda$ decreases below the elbow (rightward; toxicity weighs more), the already slow differentiation baseline rate falls toward zero, while activation/proliferation and contraction are tuned to suppress effector production, reflecting designs that minimize collateral damage even at the expense of incomplete clearance. Notably, despite being optimized only for archetypes I and II, designs in the near-elbow ensemble $\Theta(\lambda^{*})$ perform competitively for archetype III, incurring harm close to the archetype-specific optimal designs (Fig. 4E; SI Appendix, Fig. S6).
@@
Figure 6: Engineering immune response for cancer immunotherapy. (A) The schematic shows the potential clearance of replicating cancer tumor cells by T-cells (e.g., infused during immunotherapy), and the potential collateral damage of therapy to susceptible tissue cells. The accessible levers in an immunotherapy include the dosage and antigenicity of the infused cells. Antigenicity is a property of the interaction between a specific TCR-antigen pair, and it can be increased by either changing the antigen or the TCR. Given a specific cognate T-cell, immunogenicity is the ratio of the antigenicity of a cancer cell to that of a healthy cell. (B) Averaged over the tuned near-elbow ensemble $\tilde{\Theta}(\lambda^{*})$, the heatmap displays harm (in units of the maximum number of susceptible cells $S_{\text{max}})$ as a function of cancer antigenicity and dosage of administered T-cells. Yellow points correspond to the optimal dosage of T-cells for a given immunogenicity/antigenicity. (C) Total harm, harm from cancer, immune toxicity, and the final tumor size following the immune response are shown as functions of cancer immunogenicity ($\kappa$) at the optimal T-cell dosage (yellow circles in B). (D) Clearance response—defined as the derivative of negative harm (i.e., protection) with respect to each design parameter $\theta_{i}$, $\partial H(\vec{\theta})/\partial{\theta_{i}}$ for $\vec{\theta}\in\tilde{\Theta}(\lambda^{*})$—averaged over the tuned near-elbow designs $\tilde{\Theta}(\lambda^{*})$, is shown for different levels of tumor immunogenicity $\kappa$; Left: $\vec{\psi}$ parameters; right: $\vec{g}_{0}$ parameters (colors). Shaded envelopes denote the standard deviation of outcomes over $\theta\in\tilde{\Theta}(\lambda^{*})$. For this analysis, we assume that T-cells are administered at high dosage ($N_{\text{lin.}}=10^{-2}S_{\text{max}}$). Simulation parameters: $S_{\text{max}}=10^{7},K_{S}=S_{\text{max}},K_{H}=10^{-3}S_{\text{max}},d_{I}=0.0\penalty 10000\ \text{day}^{-1},I_{0}=10^{-1}S_{\text{max}},b_{I}=0.1\penalty 10000\ \text{day}^{-1}$.
@@
The choice of design impacts the dynamics of the immune response, which we can compare to experimental data. Prior experiments in mice [63, 9] indicate that peak fold-expansion of T-cell populations in response to a pathogen scales inversely with the initial number of cognate T-cells, and with the pMHC concentration for half-maximal response (proxied by $K_{I}$ in our model); see Fig. 5A, B. This pattern was recapitulated by explicitly modeling competition for antigens in polyclonal T-cell responses [66]. Without any parameter fitting, near-elbow designs $\vec{\Theta}(\lambda^{*})$ from Fig. 4 reproduce this inverse dependence on initial clone size but underestimate the steepness, and overestimate the magnitude of the trend with respect to immunogenicity ($\kappa$) (Fig. 5A, B).
Reducing the sensitivity to infection harm $\psi_{\text{inf}}$ of these designs by an order of magnitude restores the observed dynamics range (Fig. 5B) while having only a minor effect on total harm, and hence, our optimization objective (Fig. S8). Interestingly, adjusting the other sensitivities ($\psi_{\text{Ag}},\psi_{\text{res}}$), or the baseline rates $g_{0}$’s does not reproduce the dynamic range observed in the experiments (Figs. S7, S8). Improved quantitative agreement may be achieved by adjusting a subset of the model’s non-design parameters (SI Appendix, Table S1), an analysis that would benefit from a larger dataset to enable robust fitting and validation; see SI Appendix and Figs. S10–S12 for the sensitivity of model performance to variation in non-design parameters and to simulation noise.